# AI for Scientific Discovery is a Social Problem

### v2

@@ 2.1 Barrier One: Community Dysfunction
It has become popular to suggest that AGI is imminent and, that once it arrives, will solve the vast majority of our scientific problems [41]. Many efforts in the machine learning for science community have also focused on the development of so-called “AI scientists”, which have had results ranging from disappointing to actively harmful [26]. While these systems may eventually provide valuable support as co-pilot tools, particularly in retrieving accurate information and assisting with routine tasks, they are often framed as a future replacement for the human scientist.
This narrative is ultimately counterproductive to genuine scientific progress [14]. First, there is little evidence that AGI is imminent, let alone consensus on what the term means, and waiting for it risks delaying advances that could be achieved with current methods. Second, it devalues the expertise and contributions of human scientists, whose knowledge and creativity remain essential to discovery [44]. Third, it oversimplifies the inherent complexity of scientific practice, which depends not only on predictive accuracy but also on careful validation, contextualization, and theoretical integration. Finally, it obscures the central purpose of science, which is not only to produce solutions but also to cultivate human understanding [23].
@@ 2.3 Barrier Three: Data Challenges
Data fragmentation in science stems from a community that has failed to prioritize interoperability and sharing despite generating vast quantities of valuable datasets. Researchers hoard data in proprietary formats, funding agencies don’t require standardization, and institutions provide no career rewards for the tedious work of data curation and harmonization. The result is a landscape where most scientific datasets remain locked in incompatible silos, while the high-dimensional, low-sample-size nature of scientific data makes these coordination failures even more costly for AI development.
@@
The cancer research field exemplifies these challenges. Harmonizing just nine Cancer Imaging Archive files required 329.5 hours over six months, identifying only 41 overlapping fields across three or more files [basu2019call]. This labor-intensive process must be repeated for each new research initiative, creating insurmountable barriers for smaller institutions lacking dedicated data curation resources.

### v3

@@ 2.1 Barrier One: Community Dysfunction
It has become popular to suggest that artificial general intelligence (AGI) is imminent and, that once it arrives, will solve the vast majority of our scientific problems [42]. Many efforts in the machine learning for science community have also focused on the development of so-called “AI scientists”, which have had results ranging from disappointing to actively harmful [27]. While these systems may eventually provide valuable support as co-pilot tools, particularly in retrieving accurate information and assisting with routine tasks, they are often framed as a future replacement for the human scientist.
This narrative is ultimately counterproductive to genuine scientific progress [15]. First, there is little evidence that AGI is imminent, let alone consensus on what the term means, and waiting for it risks delaying advances that could be achieved with current methods. Second, it devalues the expertise and contributions of human scientists, whose knowledge and creativity remain essential to discovery [45]. Third, it oversimplifies the inherent complexity of scientific practice, which depends not only on predictive accuracy but also on careful verification, contextualization, and theoretical integration [94]. Recent work by Luo et al.  [64] demonstrates that “AI scientists” are prone to failure modes, such as inappropriate benchmark selection, data leakage, metric misuse, and post-hoc selection bias, which are likely to be overlooked when evaluating only the final output. Finally, the vision of the fully autonomous “AI scientist” obscures the central purpose of science, which is not only to produce solutions but also to cultivate human understanding [24].
@@ 2.3 Barrier Three: Data Challenges
Data fragmentation in science refers to the scattering of research data across disconnected sources, stored in incompatible formats without shared standards or effective indexing, which makes discovery, integration, and reuse difficult. It stems from a community that has failed to prioritize interoperability and sharing despite generating vast quantities of valuable datasets. Researchers hoard data in proprietary formats, funding agencies don’t require standardization, and institutions provide no career rewards for the tedious work of data curation and harmonization. The result is a landscape where most scientific datasets remain locked in incompatible silos, while the high-dimensional, low-sample-size nature of scientific data makes these coordination failures even more costly for AI development.
@@
The cancer research field exemplifies these challenges. Harmonizing just nine Cancer Imaging Archive files required 329.5 hours over six months, identifying only 41 overlapping fields across three or more files [9]. This labor-intensive process must be repeated for each new research initiative, creating insurmountable barriers for smaller institutions lacking dedicated data curation resources.